A bus travels 387 miles in 5 hours. One hour of the trip was in city traffic. The bus's city speed was just half of its speed on the open highway. The rest of the trip was on the open highway. Find the bus's city speed,highway speed, city distance, and highway distance?
Found 2 solutions by stanbon, Paul:
Answer by stanbon(75887)

City Data:
rate=x;time=1 hr.; distance =(rate)(time)=x miles
Hyway Ddata:
rate=2x; time=4 hrs.; distance = (rate)(time)=8x
EQUATION:
distance + distance = 387 miles
x+8x=387
9x=387
x=43 mph of city driving; will cover 43 miles in one hour
2x=86 mph of hyway driving; will cover 4(86)=344 miles in four hours
Cheers
stan H.

Answer by Paul(988)

One hour was in the city so 5-1 = 4 hours was in the highway.
City's speed = x/2 and highway speed = x
total distance travelled = 387

1(x/2)+4(x)=387
Multiply the equation by 2 to remove the fraction and simplfy
x+8x=774
9x=774
x=86 (highway speed)
86/2=43 (city speed)

Distacne travelled:
1(43)=43
4(86)=344

Hence, the city speed was 43mph with total distance travel in city is 43 miles and the highways speed was 86mph, with total distance on highway was 344 miles.
Paul.